OpenStudy (anonymous):

i^23 in standard a+bi form

OpenStudy (anonymous):

\[i^{23}=i^{22}\times i\]\[i ^{22}\times i =-1\times i=-i\]

OpenStudy (dumbcow):

just to clarify: \[i^{22} = (i^2)^{11} = (-1)^{11} = -1\]

OpenStudy (kainui):

since i^4=1 you can just subtract 4 until you can't anymore.